The area of the square that can be inscribed in a circle of radius 8 cm is
(A) 256 cm2 (B) 128 cm2 (C) 64 2 cm2 (D) 64 cm2

Answer by edjones(8007)

The square has a diagonal of 2*8=16
a^2+b^2=c^2
a=b
2a^2=16^2
a^2=16^2/2
=128 cm^2
.
Ed